OpenStudy (anonymous):

What is the approximate circumference of a semicircle with a radius of 12 centimeters? A. 18.8 centimeters B. 37.7 centimeters C. 75.4 centimeters D. 150.7 centimeters

OpenStudy (anonymous):

do you know whats the formula for circumference of a semicirle

OpenStudy (anonymous):

and what is pi given as

OpenStudy (anonymous):

yea, give me a second

OpenStudy (anonymous):

C=2*pi*r Semicircle C= Pi*r

OpenStudy (anonymous):

u googled it didn't you

OpenStudy (anonymous):

anyways whats the value of pi

OpenStudy (anonymous):

maybe. ^.^ 3.14?????

OpenStudy (mathmale):

Semicircle C= Pi*r: good reasoning! 3.14 is good enough as an approximation to Pi. What remains to be done?

OpenStudy (anonymous):

multiplication???????

OpenStudy (anonymous):

no its 22/7 x 12 so the answer is B

OpenStudy (anonymous):

thank u for being patient with my dumbness! ;)

OpenStudy (anonymous):

anymore questions?

OpenStudy (anonymous):

@briananicolenelson

OpenStudy (anonymous):

? confusssion!

OpenStudy (anonymous):

what confussion

OpenStudy (anonymous):

@mathmale kindly do not interfere

OpenStudy (mathmale):

@ChiefArnaz: briananicolenelson was correct in stating that the next step would be multiplication. Perhaps you did not mean to, but what you typed appears to shoot down that idea of hers. Also, 3.14 is every bit a valid approximation to Pi as is 22/7.

OpenStudy (anonymous):

....

OpenStudy (mathmale):

Briana: You were correct here: Semicircle C= Pi*r. You are told that the radius is 12 cm. You may use either 22/7 or 3.14 as an approximation to Pi: your choice. Mind calculating the circumference of the semicircle?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

so lets not argue about this...... and assume everyone's right kk?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

its your choice who u choice to know the answer

OpenStudy (anonymous):

anymore questions

OpenStudy (anonymous):

well, i dont, i am stumped...... so everyone is right yay!